Give exact and approximate solutions to three decimal places. (x-2)^2=45
Answer by stanbon(57347)

Give exact and approximate solutions to three decimal places. (x-2)^2=45
------------------------------------
x-2 = sqrt(45) or x-2 = -sqrt(45)
-----------------------------------------
x = 2+3sqrt(5) or x = 2-3sqrt(5)
============================================
Cheers,
Stan H.